# arXiv:2603.01843v1[cs.IT] 02 Mar 2026

### v1

@@ II-A Description of the TDL Channel Model
The channel is described by $N$ taps, each being the superposition of $M$ MPCs arriving with nearly the same delays. $\alpha_{n,m}$ and $\phi_{n,m}$ denote the amplitude and initial phase of the $m$-th ray belonging to tap $n$, respectively. At a given time $t$, for large $M$, by the central limit theorem $c_{n}(t)$ is well modeled by a complex Gaussian process; consequently, $|c_{n}(t)|$ is Rayleigh distributed. Empirical measurements in NLOS scenarios support the Rayleigh fading assumption.
@@ II-C Extension to MIMO Scenarios
  • Fixed Correlation Matrices: Use of $2\times 2$ correlation matrices, assumed identical across all taps for settings with $2$ transmit and receive antennas. In realistic propagation environments, this assumption is not valid, as different scatterers typically exhibit distinct spatial profiles [25].
@@ II-D Spatial Profile
In Fig. 1, we examine the spatial profile of TDL models without antenna correlation (i.e., low correlation in 3GPP parlance). We consider a MIMO system with $8$ transmit and receive antenna ports at a carrier frequency $f_{c}=3.5$ GHz, with a system bandwidth of $40$ MHz, and a sub-carrier spacing of $30$ kHz. We adopt a TDL-C profile with a delay spread of $300$ ns and a maximum Doppler frequency of $100$ Hz, corresponding to a UE speed of $30$ Km/h. This configuration is referred to as TDLC300-100 and is defined in [7, Table B.2.1.1-4].
@@
To address this issue, antenna correlations were introduced based on the Kronecker model for CSI performance requirements definition. However, the chosen antenna correlation models introduce a highly exaggerated spatial directivity towards the broadsight direction in all channel realizations, REs, and paths, which does not change over time. This behavior is illustrated in Fig. 2 for TDLC300-100 with medium correlation (spatial correlation matrix medium-A defined in [7, Sec. B.2.3.1.2]). From this observation, it also follows that TDL channel realizations employing such antenna correlation are highly rank deficient; increasing the antenna number above $3$ does not improve spatial diversity due to exceedingly small antenna correlation coefficients. Additionally, both signal and interference are consistently received from the broadsight direction, independently of the chosen precoder.
@@ III-A Description of the CDL Channel Model
The model considers $N$ clusters, each comprising $M$ rays. The channel coefficient in NLOS scenario for transmit antenna $s$, receive antenna $u$, and cluster $n$ at time $t$ is
@@ III-B Spatial Profile
A Bartlett-style spatial profile analysis, identical to that described in Sec. II-D, is applied to the CDL model. We consider the CDL-C profile (UMa scenario) with a delay spread of $365$ ns and a maximum Doppler shift of $100$ Hz in order to obtain results that are comparable with those of the TDL model. The same carrier frequency and bandwidth are used for both models. Unlike the TDL model, which defines the channel directly at the antenna port level, here we consider $128$ transmit antenna elements mapped to $8$ antenna ports as described in [3], and $8$ receive antenna ports.
@@ 1 Angular Scaling Procedure:
Here, $AS$ and $\mu_{\varphi}$ denote the angular spreads and the power weighted mean angles of the CDL-C 901 model, calculated based on the formulas provided in [1, Annex A]. Also, $c$ denotes the per-cluster spreads of arrival and departure angles given in [1, Table 7.7.1-3], and $\alpha_{m}$ represents the offset angle for ray $m$ specified in [1, Table 7.5-3]. Finally, we define
@@ 1 Type-I Codebook:
For rank $1$, the precoder is given by
@@ 1 Type-I Precoder:
We describe the rank-$1$ case; higher ranks follow similarly. After beam selection, the co-phasing value $\varphi\in\{1,j,-1,-j\}$ is chosen to minimize interference. For each candidate $\varphi$, SINRs are computed using the precoder $\mathbf{w}(\varphi)$ and channel estimates $\mathbf{H}_{j,t}$ for each physical resource block (PRB) $j\in\{0,\ldots,N_{SB}-1\}$ within SB $t$, where $N_{SB}$ is the number of PRBs per SB. Considering an LMMSE receiver, the SINR (for rank $1$) is given as
@@ 3 Effect of the Channel Estimation Accuracy:
The baseline estimation assumes equal CSI-RS and PDSCH EPRE. In practice, CSI-RS power can be boosted in the frequency domain by reallocating power from REs that are set to zero in the OFDM grid. In the following, considering pilots density and location, we apply a CSI-RS to PDSCH EPRE ratio of 4 to improve the estimation accuracy. Additionally, the DMRS bundle size is increased from $2$ to $4$ PRBs, thereby reducing estimation noise through averaging while assuming that the channel remains sufficiently stable across the bundle.

### v2

# On Channel Model to Bridge the Gap between MIMO Design and Performance Requirements in 3GPP

@@ II-A Description of the TDL Channel Model
The channel is described by NN taps, each being the superposition of $M$ MPCs arriving with nearly the same delays. $\alpha_{n,m}$ and $\phi_{n,m}$ denote the amplitude and initial phase of the $m$-th ray belonging to tap $n$, respectively. At a given time $t$, for large $M$, by the central limit theorem $c_{n}(t)$ is well modeled by a complex Gaussian process; consequently, $|c_{n}(t)|$ is Rayleigh distributed. Empirical measurements in NLOS scenarios support the Rayleigh fading assumption.
@@ II-C Extension to MIMO Scenarios
  • Fixed Correlation Matrices: Use of $2\times 2$ correlation matrices, assumed identical across all taps for settings with 22 transmit and receive antennas. In realistic propagation environments, this assumption is not valid, as different scatterers typically exhibit distinct spatial profiles [25].
@@ II-D Spatial Profile
In Fig. 1, we examine the spatial profile of TDL models without antenna correlation (i.e., low correlation in 3GPP parlance). We consider a MIMO system with 88 transmit and receive antenna ports at a carrier frequency $f_{c}=3.5$ GHz, with a system bandwidth of $40$ MHz, and a sub-carrier spacing of $30$ kHz. We adopt a TDL-C profile with a delay spread of $300$ ns and a maximum Doppler frequency of $100$ Hz, corresponding to a UE speed of $30$ Km/h. This configuration is referred to as TDLC300-100 and is defined in [7, Table B.2.1.1-4].
@@
To address this issue, antenna correlations were introduced based on the Kronecker model for CSI performance requirements definition. However, the chosen antenna correlation models introduce a highly exaggerated spatial directivity towards the broadsight direction in all channel realizations, REs, and paths, which does not change over time. This behavior is illustrated in Fig. 2 for TDLC300-100 with medium correlation (spatial correlation matrix medium-A defined in [7, Sec. B.2.3.1.2]). From this observation, it also follows that TDL channel realizations employing such antenna correlation are highly rank deficient; increasing the antenna number above 33 does not improve spatial diversity due to exceedingly small antenna correlation coefficients. Additionally, both signal and interference are consistently received from the broadsight direction, independently of the chosen precoder.
@@ III-A Description of the CDL Channel Model
The model considers NN clusters, each comprising $M$ rays. The channel coefficient in NLOS scenario for transmit antenna $s$, receive antenna $u$, and cluster $n$ at time $t$ is
@@ III-B Spatial Profile
A Bartlett-style spatial profile analysis, identical to that described in Sec. II-D, is applied to the CDL model. We consider the CDL-C profile (UMa scenario) with a delay spread of $365$ ns and a maximum Doppler shift of $100$ Hz in order to obtain results that are comparable with those of the TDL model. The same carrier frequency and bandwidth are used for both models. Unlike the TDL model, which defines the channel directly at the antenna port level, here we consider $128$ transmit antenna elements mapped to 88 antenna ports as described in [3], and 88 receive antenna ports.
@@ 1 Angular Scaling Procedure:
Here, ASAS and $\mu_{\varphi}$ denote the angular spreads and the power weighted mean angles of the CDL-C 901 model, calculated based on the formulas provided in [1, Annex A]. Also, $c$ denotes the per-cluster spreads of arrival and departure angles given in [1, Table 7.7.1-3], and $\alpha_{m}$ represents the offset angle for ray $m$ specified in [1, Table 7.5-3]. Finally, we define
@@ 1 Type-I Codebook:
For rank 11, the precoder is given by
@@ 1 Type-I Precoder:
We describe the rank-11 case; higher ranks follow similarly. After beam selection, the co-phasing value $\varphi\in\{1,j,-1,-j\}$ is chosen to minimize interference. For each candidate $\varphi$, SINRs are computed using the precoder $\mathbf{w}(\varphi)$ and channel estimates $\mathbf{H}_{j,t}$ for each physical resource block (PRB) $j\in\{0,\ldots,N_{SB}-1\}$ within SB $t$, where $N_{SB}$ is the number of PRBs per SB. Considering an LMMSE receiver, the SINR (for rank 11) is given as
@@ 3 Effect of the Channel Estimation Accuracy:
The baseline estimation assumes equal CSI-RS and PDSCH EPRE. In practice, CSI-RS power can be boosted in the frequency domain by reallocating power from REs that are set to zero in the OFDM grid. In the following, considering pilots density and location, we apply a CSI-RS to PDSCH EPRE ratio of 4 to improve the estimation accuracy. Additionally, the DMRS bundle size is increased from 22 to 44 PRBs, thereby reducing estimation noise through averaging while assuming that the channel remains sufficiently stable across the bundle.